Make New Entry,
Make Followup Entry
User name Higinbotham
Log entry time 16:12:22 on August25,2009
Entry number 285910
Hall open during the shift while Dave cooled the 25 cm target. At the end of the shift, the cooldown was almost complete. MCC called around 4pm and they are almost ready to return beam.
NOTE: The beam energy has been changed!